# Shape-specific fluctuations of an active colloidal interface

### v2

@@ VI Summary and Discussions
$W_{h}^{2}$ $\approx\frac{b^{6}}{6\tau_{f}^{2}}\langle\Big(S_{i}^{2}-\langle S\rangle^{2}\Big)^{2}\rangle t^{6}+2D_{t}t$ (30)
where we have used that $\langle F^{b}(t^{\prime})\rangle=0$ and $\langle F^{b}(t^{\prime})F^{b}(t^{\prime})\rangle=\langle F^{b}(t)\rangle\langle F^{b}(t^{\prime})\rangle$. Strictly speaking, the latter will cease to be true when curvature sets in (spring forces then become correlated), but it will suffice for the flat approximation used here. We thus obtain the result that the dynamics is super-ballistic with exponent $\sim 3$ (with an early time diffusive $\sim 0.5$). Here, fluctuations along the body axis minimally accounted for such that $\text{Var}(S)\neq 0$. We thus conclude that minimal roto-translational coupling on a flat interface can predict diffusive to super-ballistic growth transitions but clearly overestimates the super-ballistic exponent to that of a surface with finite curvature.
@@
The deterministic solution on a quasi-flat chain also enables us to rationalize the $\beta_{\theta}\sim 1$ at early times. From (27), let us probe $t<<\tau_{f}$, where we find
$\theta_{i}\approx\frac{1}{\sqrt{2}}(\chi_{r}S_{i}t)[1+\frac{\chi_{r}S_{i}t}{2}]$ (31)
@@
$\langle(\Delta\theta)^{2}\rangle$ $=\langle\theta_{i}^{2}\rangle=\frac{b^{6}}{2\tau_{f}^{2}}\langle S_{i}^{2}\rangle t^{2}+O(t^{3})+O(t^{4})$ (32)
@@
$\kappa=\left\langle\nabla\cdot\left[{\nabla h_{i}}/{\sqrt{1+|\nabla h_{i}|^{2}}}\right]\right\rangle,$ (33)
where $h_{i}$ is the height function of the chain, evaluated at each monomer location, measured from the vertical line connecting the edge monomers. The average is performed across the monomers in the chain, giving a scalar value. In our case, the inner bracket is simply reduced to $\partial_{y}\left[{\partial_{y}x_{i}(y)}/{\sqrt{1+|\partial_{y}x_{i}|^{2}}}\right]$. For Fig. 1, we take the averaged $|\kappa|$ in the steady state. Further, we note that $\kappa$ $\rightarrow 0$ as $N\rightarrow\infty$ [29]. For a circular interface, $\kappa$ would be scale as $1/R$ ($R$ the radius of the circle), whereas for a flat interface $\kappa$ would be constant. Thus, this variation the $\kappa$ profile can be used as proxy definition for the C-shape.
@@
Figure 9: Examples of DSS. From left to right: (i) Alt-C, (ii) Random front, (iii) Stiff-C, and (iv) Pinched front respectively. $\{\Lambda,\tilde{\Lambda}\}$ values are (i) $\{10^{-5},2\}$, (ii) $\{0.05,0.2\}$, (iii) $\{10^{-6},200\}$, (iv) $\{10^{1},200\}$. Note that (iv) in addition incorporated trails via (19); here $t_{0}=1$ and $\tau_{r}=5$ in addition.
@@
The scalings of $W_{\theta}$ in for $N^{\prime}<<N$ is presented in Fig. 8. We find that the scaling law of Eq. (14) is satisfied; the ballistic $\beta_{\theta}^{\prime}=1$ is retained whilst the roughness $\alpha_{\theta}^{\prime}\approx 1.02\pm 02$ is obtained. Note that this is opposed to the smoothness exponent obtained on the full C-shape. We also note that the specific choice of $N^{\prime}\in[5,21]$ used here, thus at most $\approx 4\%$ of the chain length, is substantially smaller than in Fig. 5. For any appreciably larger $N^{\prime}$ the scaling of (14) is not observed. These results thus further highlight the fact that the smoothness exponent found in Fig. 3 is a sole consequence of the C-shape; with distinct topology specific results vis-a-viz a flat surface.
@@
It is to be noted also that the phase diagram in Fig. 1 being $N$ specific and only comparing ($\Lambda$, $\tilde{\Lambda}$) does exclude other DSS that arise from simulation.Examples of these are shown in Fig. 9. Four examples are shown, we label them as (i) “Alt-C”, (ii) “Random front”, (iii) “Stiff-C”, and (iv) “Pinched front” respectively. We note that DSS(ii) corresponds to the “Disordered” phase in Fig. 1. For each of these DSS, the chain propels in a deterministic direction along the x-axis; thus (8) can be computed in principle. However, none of these DSS satisfy all the criteria enumerated in Sec. III.1. For DSS(i)-(iv), criteria (i) 2, 3, and 4, (ii) 1 and 3, (iii) 1,3 and 4; and (iv) 2 and 3 are satisfied. Thus, the C-shape is the unique topology that satisfies all four criteria. The hypothetical fluctuations for $W_{\theta}$ for these can also in principle be computed, but we do not pursue it here.
@@ Acknowledgments
We thank Professors ME Cates and M Muthukumar for useful discussions. We also thank an anonymous referee for their feedback and suggestions, which led to an improvement in the presentation of our results. AGS acknowledges funding from the DIA Fellowship from the Government of India. TB is supported through the Luxembourg National Research Fund (FNR), grant reference C22/MS/17186249.

### v3

@@ VII Appendix
$W_{h}^{2}$ $\approx\frac{b^{6}}{6\tau_{f}^{2}}\langle\Big(S_{i}^{2}-\langle S\rangle^{2}\Big)^{2}\rangle t^{6}+2D_{t}t$ (A12)
where we have used that $\langle F^{b}(t^{\prime})\rangle=0$ and $\langle F^{b}(t^{\prime})F^{b}(t^{\prime})\rangle=\langle F^{b}(t)\rangle\langle F^{b}(t^{\prime})\rangle$. Strictly speaking, the latter will cease to be true when curvature sets in (spring forces then become correlated), but it will suffice for the flat approximation used here. We thus obtain the result that the dynamics is super-ballistic with exponent $\sim 3$ (with an early time diffusive $\sim 0.5$). Here, fluctuations along the body axis are minimally accounted for such that $\text{Var}(S)\neq 0$. We thus conclude that minimal roto-translational coupling on a flat interface can predict diffusive to super-ballistic growth transitions but clearly overestimates the super-ballistic exponent to that of a surface with finite curvature.
@@
The deterministic solution on a quasi-flat chain also enables us to rationalize the $\beta_{\theta}\sim 1$ at early times. From (A9), let us probe $t<<\tau_{f}$, where we find
$\theta_{i}\approx\frac{1}{\sqrt{2}}(\chi_{r}S_{i}t)[1+\frac{\chi_{r}S_{i}t}{2}]$ (A13)
@@
$\langle(\Delta\theta)^{2}\rangle$ $=\langle\theta_{i}^{2}\rangle=\frac{b^{6}}{2\tau_{f}^{2}}\langle S_{i}^{2}\rangle t^{2}+O(t^{3})+O(t^{4})$ (A14)
@@
$\kappa=\left\langle\nabla\cdot\left[{\nabla h_{i}}/{\sqrt{1+|\nabla h_{i}|^{2}}}\right]\right\rangle,$ (A15)
where $h_{i}$ is the height function of the chain, evaluated at each monomer location, measured from the vertical line connecting the edge monomers. The average is performed across the monomers in the chain, giving a scalar value. In our case, the inner bracket is simply reduced to $\partial_{y}\left[{\partial_{y}x_{i}(y)}/{\sqrt{1+|\partial_{y}x_{i}|^{2}}}\right]$. For Fig. 1, we take the averaged $|\kappa|$ in the steady state. Further, we note that $\kappa$ $\rightarrow 0$ as $N\rightarrow\infty$ [29]. For a circular interface, $\kappa$ would be scale as $1/R$ ($R$ the radius of the circle), whereas for a flat interface $\kappa$ would be constant. Thus, this variation of the $\kappa$ profile can be used as proxy definition for the C-shape.
@@
Figure 9: Examples of DSS. From left to right: (i) Alt-C, (ii) Random front, (iii) Stiff-C, and (iv) Pinched front respectively. $\{\Lambda,\tilde{\Lambda}\}$ values are (i) $\{10^{-5},2\}$, (ii) $\{0.05,0.2\}$, (iii) $\{10^{-6},200\}$, (iv) $\{10^{1},200\}$. Note that (iv) in addition incorporated trails via (A1); here $t_{0}=1$ and $\tau_{r}=5$ in addition.
@@
The scalings of $W_{\theta}$ in for $N^{\prime}<<N$ is presented in Fig. 8. We find that the scaling law of Eq. (11) is satisfied; the ballistic $\beta_{\theta}^{\prime}=1$ is retained whilst the roughness $\alpha_{\theta}^{\prime}\approx 1.02\pm 02$ is obtained. Note that this is opposed to the smoothness exponent obtained on the full C-shape. We also note that the specific choice of $N^{\prime}\in[5,21]$ used here, thus at most $\approx 4\%$ of the chain length, is substantially smaller than in Fig. 5. For any appreciably larger $N^{\prime}$ the scaling of (11) is not observed. These results thus further highlight the fact that the smoothness exponent found in Fig. 3 is a sole consequence of the C-shape; with distinct topology-specific results vis-a-viz a flat surface.
@@
It is to be noted also that the phase diagram in Fig. 1 being $N$ specific and only comparing ($\Lambda$, $\tilde{\Lambda}$) does exclude other DSSs that arise from simulations. We show four examples in Fig. 9, where we label them as (i) “Alt-C”, (ii) “Random front”, (iii) “Stiff-C”, and (iv) “Pinched front”, respectively. We note that DSS (ii) corresponds to the “Disordered” phase in Fig. 1. For each of these DSSs, the chain propels in a deterministic direction along the x-axis; thus (6) can be computed in principle. However, none of these DSSs satisfies all the four criteria enumerated in Sec. III.1 to obtain a stable C-shape. We see that for (i) criteria 2, 3, and 4, for (ii) 1 and 3, for (iii) 1,3 and 4; and for (iv) 2 and 3 are satisfied. Thus, the C-shape is the unique topology that satisfies all the four criteria. The fluctuations for $W_{\theta}$ for each of these shapes can also, in principle, be computed, but we do not pursue such questions here.
@@ Acknowledgments
We thank Professors ME Cates and M Muthukumar for useful discussions. We also thank the anonymous referee for their feedback and suggestions, which led to an improvement in the presentation of our results. AGS acknowledges funding from the DIA Fellowship from the Government of India. TB is supported through the Luxembourg National Research Fund (FNR), grant reference C22/MS/17186249.